# Training Large Language Models to Reason in a Continuous Latent Space

### v2

@@ 4.4 Results and Discussion
Table 1: Results on three datasets: GSM8l, ProntoQA and ProsQA. Higher accuracy indicates stronger reasoning ability, while generating fewer tokens indicates better efficiency. The result is from Deng et al. (2024).
Figure 3: Accuracy on GSM8k with different number of continuous thoughts.
We show the overall results on all datasets in Table 1. Continuous thoughts effectively enhance LLM reasoning, as shown by the consistent improvement over no-CoT. It even shows better performance than CoT on ProntoQA and ProsQA. We describe several key conclusions from the experiments as follows.
“Chaining” continuous thoughts enhances reasoning. In conventional CoT, the output token serves as the next input, which proves to increase the effective depth of LLMs and enhance their expressiveness (Feng et al., 2023). We explore whether latent space reasoning retains this property, as it would suggest that this method could scale to solve increasingly complex problems by chaining multiple latent thoughts.
In our experiments with GSM8k, we found that Coconut outperformed other architectures trained with similar strategies, particularly surpassing the latest baseline, iCoT (Deng et al., 2024). The performance is significantly better than Coconut (pause as thought) which also enables more computation in the LLMs. While Pfau et al. (2024) empirically shows that filler tokens, such as the special <pause> tokens, can benefit highly parallelizable problems, our results show that Coconut architecture is more effective for general problems, e.g., math word problems, where a reasoning step often heavily depends on previous steps. Additionally, we experimented with adjusting the hyperparameter $c$, which controls the number of latent thoughts corresponding to one language reasoning step (Figure 3). As we increased $c$ from 0 to 1 to 2, the model’s performance steadily improved. [^2] We discuss the case of larger $c$ in Appendix 9. These results suggest that a chaining effect similar to CoT can be observed in the latent space.
@@
In two other synthetic tasks, we found that the variants of Coconut (w/o thoughts or pause as thought), and the iCoT baseline also achieve impressive accuracy. This indicates that the model’s computational capacity may not be the bottleneck in these tasks. In contrast, GSM8k, being an open-domain question-answering task, likely involves more complex contextual understanding and modeling, placing higher demands on computational capability.
Latent reasoning outperforms language reasoning in planning-intensive tasks. Complex reasoning often requires the model to “look ahead” and evaluate the appropriateness of each step. Among our datasets, GSM8k and ProntoQA are relatively straightforward for next-step prediction, due to intuitive problem structures and limited branching. In contrast, ProsQA’s randomly generated DAG structure significantly challenges the model’s planning capabilities. As shown in Table 1, CoT does not offer notable improvement over No-CoT. However, Coconut, its variants, and iCoT substantially enhance reasoning on ProsQA, indicating that latent space reasoning provides a clear advantage in tasks demanding extensive planning. An in-depth analysis of this process is provided in Section 5.
@@
Figure 4: A case study where we decode the continuous thought into language tokens.
@@
With the multi-stage curriculum which decomposes the training into easier objectives, Coconut is able to achieve top performance across various tasks. The multi-stage training also integrates well with pause tokens (Coconut- pause as thought). Despite using the same architecture and similar multi-stage training objectives, we observed a small gap between the performance of iCoT and Coconut (w/o thoughts). The finer-grained removal schedule (token by token) and a few other tricks in iCoT may ease the training process. We leave combining iCoT and Coconut as future work. While the multi-stage training used for Coconut has proven effective, further research is definitely needed to develop better and more general strategies for learning reasoning in latent space, especially without the supervision from language reasoning chains.
Continuous thoughts are efficient representations of reasoning. Though the continuous thoughts are not intended to be decoded to language tokens, we can still use it as an intuitive interpretation of the continuous thought. We show a case study in Figure 4 of a math word problem solved by Coconut ($c=1$). The first continuous thought can be decoded into tokens like “180”, “ 180” (with a space), and “9”. Note that, the reasoning trace for this problem should be $3\times 3\times 60=9\times 60=540$, or $3\times 3\times 60=3\times 180=540$. The interpretations of the first thought happen to be the first intermediate variables in the calculation. Moreover, it encodes a distribution of different traces into the continuous thoughts. As shown in Section 5.3, this feature enables a more advanced reasoning pattern for planning-intense reasoning tasks.
5 Understanding the Latent Reasoning in Coconut
In this section, we present an analysis of the latent reasoning process with a variant of Coconut. By leveraging its ability to switch between language and latent space reasoning, we are able to control the model to interpolate between fully latent reasoning and fully language reasoning and test their performance (Section 5.2). This also enables us to interpret the the latent reasoning process as tree search (Section 5.3). Based on this perspective, we explain why latent reasoning can make the decision easier for LLMs (Section 5.4).
5.1 Experimental Setup
Methods. The design of Coconut allows us to control the number of latent thoughts by manually setting the position of the <eot> token during inference. When we enforce Coconut to use $k$ continuous thoughts, the model is expected to output the remaining reasoning chain in language, starting from the $k+1$ step. In our experiments, we test variants of Coconut on ProsQA with $k\in\{0,1,2,3,4,5,6\}$. Note that all these variants only differ in inference time while sharing the same model weights. Besides, we report the performance of CoT and no-CoT as references.
To address the issue of forgetting earlier training stages, we modify the original multi-stage training curriculum by always mixing data from other stages with a certain probability ($p=0.3$). This updated training curriculum yields similar performance and enables effective control over the switch between latent and language reasoning.
Metrics. We apply two sets of evaluation metrics. One of them is based on the correctness of the final answer, regardless of the reasoning process. It is the metric used in the main experimental results above (Section 1). To enable fine-grained analysis, we define another metric on the reasoning process. Assuming we have a complete language reasoning chain which specifies a path in the graph, we can classify it into (1) Correct Path: The output is one of the shortest paths to the correct answer. (2) Longer Path: A valid path that correctly answers the question but is longer than the shortest path. (3) Hallucination: The path includes nonexistent edges or is disconnected. (4) Wrong Target: A valid path in the graph, but the destination node is not the one being asked. These four categories naturally apply to the output from Coconut ($k=0$) and CoT, which generate the full path. For Coconut with $k>0$ that outputs only partial paths in language (with the initial steps in continuous reasoning), we classify the reasoning as a Correct Path if a valid explanation can complete it. Also, we define Longer Path and Wrong Target for partial paths similarly. If no valid explanation completes the path, it’s classified as hallucination. In no-CoT and Coconut with larger $k$, the model may only output the final answer without any partial path, and it falls into (5) Correct Label or (6) Incorrect Label. These six categories cover all cases without overlap.
5.2 Interpolating between Latent and Language Reasoning
Figure 5 shows a comparative analysis of different reasoning methods on ProsQA. As more reasoning is done with continuous thoughts (increasing $k$), both final answer accuracy (Figure 5, left) and the rate of correct reasoning processes (“Correct Label” and “Correct Path” in Figure 5, right) improve. Additionally, the rate of “Hallucination” and “Wrong Target” decrease, which typically occur when the model makes a wrong move earlier. This also indicates the better planning ability when more reasoning happens in the latent space.
A case study is shown in Figure 7, where CoT hallucinates an nonexistent edge, Coconut ($k=1$) leads to a wrong target, but Coconut ($k=2$) successfully solves the problem. In this example, the model cannot accurately determine which edge to choose at the earlier step. However, as latent reasoning can avoid making a hard choice upfront, the model can progressively eliminate incorrect options in subsequent steps and achieves higher accuracy at the end of reasoning. We show more evidence and details of this reasoning process in Section 5.3.
The comparison between CoT and Coconut ($k=0$) reveals another interesting observation: even when Coconut is forced to generate a complete reasoning chain, the accuracy of the answers is still higher than CoT. The generated reasoning paths are also more accurate with less hallucination. From this, we can infer that the training method of mixing different stages improves the model’s ability to plan ahead. The training objective of CoT always concentrates on the generation of the immediate next step, making the model “shortsighted”. In later stages of Coconut training, the first few steps are hidden, allowing the model to focus more on future steps. This is related to the findings of Gloeckle et al. (2024), where they propose multi-token prediction as a new pretraining objective to improve the LLM’s ability to plan ahead.
5.3 Interpreting the Latent Search Tree
Given the intuition that continuous thoughts can encode multiple potential next steps, the latent reasoning can be interpreted as a search tree, rather than merely a reasoning “chain”. Taking the case of Figure 7 as a concrete example, the first step could be selecting one of the children of Alex, i.e., {lempus, sterpus, zhorpus, grimpus}. We depict all possible branches in the left part of Figure 7. Similarly, in the second step, the frontier nodes will be the grandchildren of Alex (Figure 7, right).
Unlike a standard breadth-first search (BFS), which explores all frontier nodes uniformly, the model demonstrates the ability to prioritize promising nodes while pruning less relevant ones. To uncover the model’s preferences, we analyze its subsequent outputs in language space. For instance, if the model is forced to switch back to language space after a single latent thought ($k=1$), it predicts the next step in a structured format, such as “every [Concept A] is a [Concept B].” By examining the probability distribution over potential fillers for [Concept A], we can derive numeric values for the children of the root node Alex (Figure 7, left). Similarly, when $k=2$, the prediction probabilities for all frontier nodes—the grandchildren of Alex—are obtained (Figure 7, right).
Figure 8: Analysis of parallelism in latent tree search. The left plot depicts the cumulative value of the top-1, top-2, and top-3 candidate nodes for the first thoughts, calculated across test cases and ranked by percentile. The significant gaps between the lines reflect the model’s ability to explore alternative latent thoughts in parallel. The right plot shows the corresponding analysis for the second thoughts, where the gaps between lines are narrower, indicating reduced parallelism and increased certainty in reasoning as the search tree develops. This shift highlights the model’s transition toward more focused exploration in later stages.
The probability distribution can be viewed as the model’s implicit value function, estimating each node’s potential to reach the target. As shown in the figure, “lempus”, “zhorpus”, “grimpus”, and “sterpus” have a value of 0.33, 0.16, 0.32, and 0.01, respectively. This indicates that in the first continuous thought, the model has mostly ruled out “sterpus” as an option but remains uncertain about the correct choice among the other three. In the second thought, however, the model has mostly ruled out other options but focused on “rorpus”.
5.4 Why is a Latent Space Better for Planning?
Figure 9: The correlation between prediction probability of concepts and their heights.
In this section, we explore why latent reasoning is advantageous for planning, drawing on the search tree perspective and the value function defined earlier. Referring to our illustrative example, a key distinction between “sterpus” and the other three options lies in the structure of the search tree: “sterpus” is a leaf node (Figure 7). This makes it immediately identifiable as an incorrect choice, as it cannot lead to the target node “bompus”. In contrast, the other nodes have more descendants to explore, making their evaluation more challenging.
To quantify a node’s exploratory potential, we measure its height in the tree, defined as the shortest distance to any leaf node. Based on this notion, we hypothesize that nodes with lower heights are easier to evaluate accurately, as their exploratory potential is limited. Consistent with this hypothesis, in our example, the model exhibits greater uncertainty between “grimpus” and “lempus”, both of which have a height of 2—higher than the other candidates.
To test this hypothesis more rigorously, we analyze the correlation between the model’s prediction probabilities and node heights during the first and second latent reasoning steps across the test set. Figure 9 reveals a clear pattern: the model successfully assigns lower values to incorrect nodes and higher values to correct nodes when their heights are low. However, as node heights increase, this distinction becomes less pronounced, indicating greater difficulty in accurate evaluation.
In conclusion, these findings highlight the benefits of leveraging latent space for planning. By delaying definite decisions and expanding the latent reasoning process, the model pushes its exploration closer to the search tree’s terminal states, making it easier to distinguish correct nodes from incorrect ones.
@@ 6 Conclusion
In this paper, we presented Coconut, a novel paradigm for reasoning in continuous latent space. Through extensive experiments, we demonstrated that Coconut significantly enhances LLM reasoning capabilities. Notably, our detailed analysis highlighted how an unconstrained latent space allows the model to develop an effective reasoning pattern similar to BFS. Future work is needed to further refine and scale latent reasoning methods. One promising direction is pretraining LLMs with continuous thoughts, which may enable models to generalize more effectively across a wider range of reasoning scenarios. We anticipate that our findings will inspire further research into latent reasoning methods, ultimately contributing to the development of more advanced machine reasoning systems.

### v3

@@ 5 Empirical Results with Coconut
Table 1: Results on three datasets: GSM8k, ProntoQA and ProsQA. Higher accuracy indicates stronger reasoning ability, while generating fewer tokens indicates better efficiency. The result is from Deng et al. (2024).
5.1 Experimental Setup
Math Reasoning. We use GSM8k (Cobbe et al., 2021) as the dataset for math reasoning. It consists of grade school-level math problems. To train the model, we use a synthetic dataset generated by Deng et al. (2023). We use two continuous thoughts for each reasoning step (i.e., $c=2$). The model goes through 3 stages besides the initial stage. We then include an additional stage where still $3\times c$ continuous thoughts are used as in the previous stage, but with all the remaining language reasoning chain removed. This handles the long-tail distribution of reasoning chains longer than 3 steps. We train the model for 6 epochs in the initial stage, and 3 epochs in each remaining stage.
Logical Reasoning. Logical reasoning involves the proper application of known conditions to prove or disprove a conclusion using logical rules. We use the ProntoQA (Saparov and He, 2022) dataset, and our newly proposed ProsQA dataset, which is more challenging due to more distracting branches. We use one continuous thought for every reasoning step (i.e., $c=1$). The model goes through 6 training stages in addition to the initial stage, because the maximum number of reasoning steps is 6 in these two datasets. The model then fully reasons with continuous thoughts to solve the problems in the last stage. We train the model for 5 epochs per stage.
For all datasets, after the standard schedule, the model stays in the final training stage, until reaching 50 epochs. We select the checkpoint based on the accuracy on the validation set. For inference, we manually set the number of continuous thoughts to be consistent with their final training stage. We use greedy decoding for all experiments.
5.2 Baselines and Variants of Coconut
We consider the following baselines: (1) CoT, and (2) No-CoT, which were introduced in Section 4. (3) iCoT (Deng et al., 2024): The model is trained with language reasoning chains and follows a carefully designed schedule that “internalizes” CoT. As the training goes on, tokens at the beginning of the reasoning chain are gradually removed until only the answer remains. During inference, the model directly predicts the answer. (4) Pause token (Goyal et al., 2023): The model is trained using only the question and answer without a reasoning chain. However, different from No-CoT, special <pause> tokens are inserted between the question and answer, which provides the model with additional computational capacity to derive the answer. The number of <pause> tokens is set the same as continuous thoughts in Coconut.
We also evaluate some variants of Coconut: (1) w/o curriculum, which directly trains the model in the last stage. The model uses continuous thoughts to solve the whole problem. (2) w/o thought: We keep the multi-stage training, but don’t add any continuous latent thoughts. While this is similar to iCoT in the high-level idea, the exact training schedule is set to be consistent with Coconut, instead of iCoT, for a strict comparison. (3) Pause as thought: We use special <pause> tokens to replace the continuous thoughts, and apply the same multi-stage training curriculum as Coconut.
5.3 Results and Discussion
We show the overall results in Table 1. Using continuous thoughts effectively enhances LLM reasoning over the No-CoT baseline. For example, by using 6 continuous thoughts, Coconut achieves 34.1% accuracy on GSM8k, which significantly outperforms No-CoT (16.5%). We highlight several key findings below.
Figure 8: Efficiency comparison of reasoning space and Coconut with different $c$.
“Chaining” continuous thoughts enhances reasoning. Language CoT proves to increase the effective depth of LLMs and enhance their expressiveness (Feng et al., 2023). Thus, generating more tokens serves as a way to inference-time scaling for reasoning (Guo et al., 2025; Snell et al., 2024). This desirable property holds naturally for Coconut too. On GSM8k, Coconut outperformed other architectures trained with similar strategies, including Coconut (pause as thought) and Coconut (w/o thought). Particularly, it surpasses the latest baseline iCoT (Deng et al., 2024), which requires a more carefully designed training schedule.
@@ 5.3 Results and Discussion
Additionally, we experimented with adjusting the hyperparameter $c$, which controls the number of latent thoughts corresponding to one language reasoning step (Figure 8, II). As we increased $c$ from 0 to 1 to 2, the model’s performance steadily improved. [^2] We discuss the case of larger $c$ in Appendix 9.1. This further validates the potential of continuous thoughts to scale up to harder problems. In two other synthetic tasks, we found that the variants of Coconut (w/o thoughts or pause as thought), and the iCoT baseline also achieve impressive accuracy. This indicates that the model’s computational capacity may not be the bottleneck in these tasks. In contrast, GSM8k involves more complex contextual understanding and modeling, placing higher demands on computational capability.
Figure 9: Decoding a continuous thought into language tokens in a math word problem. The decoded tokens correspond to intermediate variables that help solve the problem.
Continuous thoughts are efficient representations of reasoning. Compared to traditional CoT, Coconut generates fewer tokens while achieving higher accuracy on ProntoQA and ProsQA (Table 1). Although Coconut does not surpass CoT on GSM8k, it offers a superior trade-off between reasoning efficiency and accuracy (Figure 8, I). To illustrate this, we train a series of CoT models that progressively “internalize” (Deng et al., 2024) the initial $m=\{0,1,2,3,\text{ALL}\}$ reasoning steps, and plot their accuracy versus the number of generated tokens (labeled as “language” in the figure). These models quickly lose accuracy as more reasoning steps are skipped. In contrast, by applying Coconut training strategy—replacing each language reasoning step with two continuous thoughts—the accuracy drop is substantially mitigated, maintaining higher performance even when fewer tokens are generated. Another interesting observation is that, when we decode the first continuous thought, it often corresponds to possible intermediate variables in the calculation (Figure 9). This also suggests that the continuous thoughts are more efficient representations of reasoning.
@@
On the contrary, with the multi-stage curriculum, Coconut is able to achieve top performance across various tasks. The multi-stage training also integrates well with pause tokens (Coconut- pause as thought). Despite using the same architecture and similar multi-stage training objectives, we observed a small gap between the performance of iCoT and Coconut (w/o thoughts). The finer-grained removal schedule (token by token) and a few other tricks in iCoT may ease the training process. We leave combining iCoT and Coconut as future work. While the multi-stage training used for Coconut has proven effective, further research is definitely needed to develop better and more general strategies for learning reasoning in latent space, especially without the supervision from language reasoning chains.
@@ 6 Conclusion
In this paper, we introduce Coconut, a new paradigm for reasoning in continuous latent space. Experiments demonstrate that Coconut effectively enhances LLM performance across a variety of reasoning tasks. Reasoning in latent space gives rise to advanced emergent behaviors, where continuous thoughts can represent multiple alternative next steps. This enables the model to perform BFS over possible reasoning paths, rather than prematurely committing to a single deterministic trajectory as in language space CoT reasoning. Further research is needed to refine and scale latent reasoning to pretraining, which could improve generalization across a broader range of reasoning challenges. We hope our findings will spark continued exploration into latent reasoning, ultimately advancing the development of more capable machine reasoning systems.